### v1

# DAG-Math: Graph-Guided Mathematical Reasoning in LLMs

Abstract
Large Language Models (LLMs) demonstrate strong performance on mathematical problems when prompted with Chain-of-Thought (CoT), yet it remains unclear whether this success stems from search, rote procedures, or rule-consistent reasoning. To address this, we propose modeling CoT as a certain rule-based stochastic process over directed acyclic graphs (DAGs), where nodes represent intermediate derivation states and edges encode rule applications. Within this framework, we introduce logical closeness, a metric that quantifies how well a model’s CoT trajectory (i.e., the LLM’s final output) adheres to the DAG structure, providing evaluation beyond classical PASS@$k$ metrics. Building on this, we introduce the DAG-MATH CoT format and construct a benchmark that guides LLMs to generate CoT trajectories in this format, thereby enabling the evaluation of their reasoning ability under our framework. Across standard mathematical reasoning datasets, our analysis uncovers statistically significant differences in reasoning fidelity among representative LLM families—even when PASS@$k$ is comparable—highlighting gaps between final-answer accuracy and rule-consistent derivation. Our framework provides a balance between free-form CoT and formal proofs systems, offering actionable diagnostics for LLMs reasoning evaluation. Our benchmark and code are available at https://github.com/YuanheZ/DAG-MATH-Formatted-CoT.
1 Introduction
Large Language Models (LLMs) have demonstrated promising mathematical reasoning abilities on answer/proof-based problems, e.g., Gemini-2.5 (Gemini Team, 2025) and GPT-5 (OpenAI, 2025), DeepSeek-R1 (DeepSeek Team, 2025). A key strategy underlying these successes is the Chain-of-Thought (CoT) (Nye et al., 2021; Wei et al., 2022; Kojima et al., 2022; Zhang et al., 2022), which encourages models to produce intermediate reasoning steps prior to the final answer.
The black-box nature of CoT in LLMs raises a key challenge: how to rigorously model and evaluate LLMs’ mathematical reasoning abilities. Prior works formalize reasoning in terms of search (Shalev-Shwartz & Shashua, 2025), probabilistic inference (Prystawski et al., 2023; Feng et al., 2023), and propositional logic (Merrill & Sabharwal, 2023; Kim & Suzuki, 2024; Yin et al., 2025). Intuitively, LLM reasoning needs to identify all required premises (e.g., facts, constraints), and conduct correct logical inference from premises to reach the conclusion. [^1] Accurate calculation and symbolic execution are also required, see the discussion in Section A.1. These operations must be exact, in line with the exact learning requirements in György et al. (2025). To test whether LLMs achieve this through CoT, two elements are crucial:
  • A Rigorous Framework: It is necessary to establish a principled framework that characterizes the mechanisms by which CoT operates in mathematical problem solving. Such a framework should explicitly account for the roles of premises identification and logical inference. This in turn enables a systematic analysis of reasoning behaviors.
  • An Appropriate Evaluation Metric: A reliable metric is required to assess whether model outputs reflect authentic reasoning processes rather than the application of heuristic or search-based strategies [^2] Search-based strategies may yield irrelevant information, undermining solution’s consistency. LLMs should be able to summarize the searched/thinking results to ensure the final output logic coherence.. An effective metric should hence evaluate not only the final correctness but also the logical coherence of intermediate reasoning steps.
Despite recent progress, existing approaches remain limited in both framework and evaluation. In terms of frameworks, prior work has examined Boolean-circuit analyses (Li et al., 2024), $k$-parity models (Kim & Suzuki, 2024; Yin et al., 2025), and graph-based representations (Dziri et al., 2023). Graph-based strategies provide a natural way to model CoT as a discrete, step-level, graph-based abstractions where nodes correspond to intermediate assertions and edges encode inferential dependencies. However, existing graph-based formulations, whether modeling CoT as deterministic subgraph matching in a directed acyclic graph (DAG) (Dziri et al., 2023), random walks on reversible/recurrent Markov chains (Kim et al., 2025), or trees of linear solution paths (Shalev-Shwartz & Shashua, 2025), neglect the improvement from diverse sampling (Wang et al., 2023) and ability to connect disparate knowledge (Yin & Wang, 2025). As a result, they fail to capture long-range and cross-branch dependencies, as well as the goal-directed, absorbing-state nature of CoT.
In terms of evaluations, prior work (Dziri et al., 2023; Joshi et al., 2025; Kim et al., 2025; Xu & Sato, 2025) primarily relies on final-answer metrics such as PASS@$k$ while the entire output for inference is overlooked, which leaves it unclear whether it is solved by logical inference or by search. A promising alternative is to use the LEAN programming language (De Moura et al., 2015; Moura & Ullrich, 2021) to formally verify solutions (Google DeepMind, 2024; Ren et al., 2025; Wang et al., 2025; Lin et al., 2025). While LEAN-based verification ensures logical correctness, it presupposes that problems have been already formalized in a proof-oriented form. For answer-based problems, the formalization into Lean must be carried out in advance, which requires substantial expert effort.
Building on empirical and theoretical insights (Ye et al., 2025; Kim et al., 2025), we address these limitations by formalizing CoT as a rule-based stochastic process on directed acyclic graphs (DAGs). This formalization provides a unified and principled framework for both modeling and evaluation of LLM mathematical reasoning. The main contributions of this work are summarized as below:
  • Framework. In Section˜2, we establish a rigorous graph-based framework that formalizes CoT (i.e., the LLM’s entire output) in two phases. Phase 1 constructs a task-specific DAG as the search space for generating CoT trajectories. In Phase 2, the LLM generates CoT trajectories over this DAG under certain stochastic transition rules.
  • Logical Closeness and Metric. In Section˜3, we introduce the notion of logical closeness, which evaluates whether an LLM solves a problem by searching over possible choices or by applying rigorous logical inference. This yields a new evaluation metric, the perfect reasoning rate (PRR) and the related AUC (area under curve) scores for ranking.
  • Benchmark Construction. In Section˜4, we propose the DAG-MATH format based on the above concepts, which makes the logical structure of CoT explicit through DAG representations. Using a three-stage prompting method, we construct a benchmark of 2,894 gold-standard DAG-MATH DAGs. Our graph-level statistical analysis shows that harder problems yield larger, sparser DAGs with higher branching complexity, emphasizing the need for LLMs to decompose tasks, track long dependencies, and recombine results effectively.
  • Empirical Evaluation. In Section˜5, we employ few-shot prompting to guide LLMs (e.g., Gemini, GPT, Qwen3) to produce DAG-MATH formatted CoT trajectories as the final output. We find that search can inflate PASS@1 through exploratory branching, while perfect reasoning ability remains comparable across models. Perfectly reasoning corresponds to easier problems; only-correct-answer CoT reflects modest exploratory overhead; and incorrect CoT typically arises from problems exceeding model capacity, where difficulty stems from branching rather than aggregation.
Our framework provides a “Goldilocks principle” that balances the versatility of natural language with the rigor of LEAN in mathematical reasoning evaluation. Moreover, we believe the DAG-MATH framework can lay the foundation for a mathematical definition of reasoning in LLMs (paralleling in memorization and generalization in supervised learning), and inform algorithm design for improved reasoning performance of LLMs, leaving for further investigation.
Notations: We denote a random variable by a capital letter (e.g., $V$) and its realization by the corresponding lowercase letter (e.g., $v$). For shorthand, we write $v_{1:t}=(v_{1}\,,v_{2}\,,\ldots\,,v_{t})$ for $t\geq 1$. We denote a DAG by ${\mathcal{G}}=({\mathcal{E}}\,,{\mathcal{V}})$, where $\mathcal{V}$} is the node set and $\mathcal{E}$} is the edge set. For a node $v$, we write $\operatorname{pa}(v)$ as its parent set. Finally, we denote the input prompt by ${\bm{X}}_{\tt in}\in{\mathcal{P}}$, where $\mathcal{P}$} is the power set of the vocabulary.
2 A DAG Framework for Step-Level CoT
Motivated by empirical observations in Bogdan et al. (2025), we study CoT at the step level, rather than the token level. This step-level perspective has been widely considered in recent theoretical analyses (Dziri et al., 2023; Hu et al., 2024; Kim et al., 2025; Shalev-Shwartz & Shashua, 2025), as it better captures intermediate reasoning and the logical structure of solutions. We model step-level CoT in a two-phase workflow as below. Phase 1 defines a task-specific DAG, where Phase 2 samples CoT trajectories over this DAG under certain stochastic process. For better illustration, we take the following problem, adapted from MATH-500 (Hendrycks et al., 2021), as a representative example.
Logarithmic Count Problem (LCP) For how many integers $k\in[-300,300]$ does the equation $2\log(x-1)=\log k$ have exactly one real solution $x$?
2.1 Phase 1: Task-specific DAG for Step-Level CoT
Edges and Nodes in Step-Level CoT:
For mathematical problems, a CoT step is a natural-language derivation of a new conclusion from prior information. Each step has two components:
  • Edge (Justification):This captures the inference that leads to the step’s conclusion. The edge explicitly encodes the logical dependency on the problem statement or on previous steps, making the reasoning chain transparent.
  • Node (Conclusion): The node represents the step’s conclusion—the state or value derived from the edge’s logic and its parent nodes.
Hence, a single CoT step can be viewed as node/edge decomposition, see an example below as well as the nodes/edges in Fig.˜1 for the logarithmic count problem.
One CoT Step in the Logarithmic Count Problem (LCP) If two logarithmic expressions are equal, then their arguments must also be equal. Hence, from $2\log(x-1)=\log k$, we can conclude $(x-1)^{2}=k$.
Here the blue part corresponds to the previous conclusion (parent node), the green part represents the new conclusion for the current node, and the orange part highlights the logical reasoning (Edge) that connects the parent to the current node. Note that the edge may be latent when the model only outputs the conclusion without explicitly stating the reasoning.
Note that such a decomposition is non-unique due to semantic variation, such as synonyms or equivalent phrasings. This ambiguity makes it challenging to develop a precise and principled definition of a CoT step. Consequently, prior work (Dziri et al., 2023; Hu et al., 2024; Kim et al., 2025; Shalev-Shwartz & Shashua, 2025; Bogdan et al., 2025) has typically defined steps heuristically—either as text spans or task-specific annotations—without providing a formal definition. As a first attempt, we present an abstract mathematical formulation, with the technical details deferred to Section˜A.2 since they are not essential for understanding the main text. Within this formulation, although the node/edge decomposition for a single step may still be non-unique, the task-specific DAG introduced later can be made unique, provided that each step is restricted to a single conclusion.
Task-Specific DAG:
Empirical studies (Ye et al., 2025) demonstrate the existence of a latent directed dependency graph within LLMs, present as soon as a question/prompt is posted, before any output is generated. Formally, given a prompt ${\bm{x}}_{\tt in}$, we define the directed graph as
${\mathcal{G}}({\bm{x}}_{\tt in}):=({\mathcal{V}}({\bm{x}}_{\tt in})\,,{\mathcal{E}}({\bm{x}}_{\tt in}))\,,\quad\mbox{where}~\,{\mathcal{E}}({\bm{x}}_{\tt in})\subseteq{\mathcal{V}}({\bm{x}}_{\tt in})\times{\mathcal{V}}({\bm{x}}_{\tt in})\,$,
where ${\mathcal{E}}({\bm{x}}_{\tt in})$ is the set of directed edges and ${\mathcal{V}}({\bm{x}}_{\tt in})$ is the set of nodes divided into three classes:
  • ${\mathcal{V}}_{\tt in}({\bm{x}}_{\tt in})$ denotes the set of source nodes, i.e., nodes formulated solely from the input prompt. In Fig.˜1, the source nodes are $v_{1},v_{2},$ and $v_{3}$.
  • ${\mathcal{V}}_{\tt out}({\bm{x}}_{\tt in})$ denotes the set of sink nodes, i.e., nodes with only incoming edges and no outgoing edges, corresponding to the final answer(s). The correct sink node represents the terminal object that matches the ground-truth answer. In Fig.˜1, the sink nodes are $v_{10}$ (correct) and $v_{11}$ (incorrect).
  • ${\mathcal{V}}_{\tt inter}({\bm{x}}_{\tt in}):={\mathcal{V}}({\bm{x}}_{\tt in})\setminus\bigl({\mathcal{V}}_{\tt in}({\bm{x}}_{\tt in})\cup{\mathcal{V}}_{\tt out}({\bm{x}}_{\tt in})\bigr)$ denotes the set of intermediate nodes. In Fig.˜1, the intermediate nodes are $v_{4}$ through $v_{9}$.
We make the following assumption on the acyclic structure of the graph for the absence of circular dependencies, ensuring that no CoT step depends on its own output either directly or indirectly.
Assumption 2.1.
For any input prompt ${\bm{x}}_{\tt in}$, the task-specific directed graph ${\mathcal{G}}({\bm{x}}_{\tt in})$ is acyclic.
This assumption covers the answer-based math problems which have tractable computation graphs such as AIME (Art of Problem Solving, 2025a, b). Note that, if the correct sink node is included in ${\mathcal{G}}({\bm{x}}_{\tt in})$, the task-specific DAG can be always constructed by backtracking through its ancestors. We remark that, some “thinking-LLMs”, e.g., DeepSeek-R1 (DeepSeek Team, 2025), encourage backtracking and self-correction during the thinking process. However, we expect that LLMs output only the finalized, perfect, correct reasoning results to the user. The reasoning evaluation in this paper is based on the entire final output (while PASS@$k$ just considers the final answer). The definition of perfect and correct reasoning will be introduced in the next section.
Figure 1: Task-specific DAG via LLaMA-3.1-8B-Instruct.
[…]
  • $v_{1}:$$2\log(x-1)=\log k$" (target equation);
  • $v_{2}:$$k\in[-300\,,300]$" (range constraint);
  • $v_{3}:$ “exactly one solution" (task requirement);
  • $v_{4}:$$\log(x-1)$ requires x>1x>1" (constraint inferred from the equation);
  • $v_{5}:$$\log k$ requires k>0k>0" (constraint inferred from the equation);
  • $v_{6}:$$2\log(x-1)=\log k\;\Rightarrow\;(x-1)^{2}=k$" (re-arranged equation);
  • $v_{7}:$$x=1\pm\sqrt{k}$" (solve the quadratic equation);
  • $v_{8}:$$1+\sqrt{k}$ is the only solution" (correct check);
  • $v_{9}:$ “For any $k$, there are two solutions" (incorrect check);
  • $v_{10}:$ “There are $300$ valid values for kk" (the correct answer);
  • $v_{11}:$ “There are 0 valid value for kk" (the incorrect answer).
2.2 Phase 2: Stochastic Process on Logic Dependence
Based on the task-specific DAG, the LLM generates CoT trajectories over this DAG as the final output via a certain sampling strategy. Given ${\mathcal{G}}({\bm{x}}_{\tt in})$ from Phase 1, we denote the node-level autoregressive distribution of an LLM as $\mathbb{P}$}. A node-level CoT trajectory $\{V_{i}\}_{i=1}^{L}$ with length-$L$, given the input prompt ${\bm{X}}_{\tt in}$, sequentially generates $V_{t}\in{\mathcal{V}}$ ($1\leq t\leq L$), ultimately leading to the final answer $V_{L}:=V_{\tt out}$. Specifically, the trajectory $\{V_{i}\}_{i=1}^{L}$ follows the stochastic process:
$V_{1}\!\sim\!{\mathbb{P}}(\,\cdot\mid{\bm{X}}_{\tt in})\,,\cdots\,,V_{t}\sim{\mathbb{P}}(\,\cdot\mid V_{t-1}\,,\ldots\,,V_{1}\,,{\bm{X}}_{\tt in})\,,\cdots,~V_{\tt out}\sim{\mathbb{P}}(\,\cdot\mid V_{L-1}\,,\cdots\,,V_{1}\,,{\bm{X}}_{\tt in})\,.$
Next, we define a stochastic transition rule to generate the node-level trajectory over ${\mathcal{V}}({\bm{x}}_{\tt in})$. We begin with the initial step, where ${\mathbb{P}}(V_{1}=v\mid{\bm{X}}_{\tt in}={\bm{x}}_{\tt in})$ is nonzero only for $v\in{\mathcal{V}}_{\tt in}({\bm{x}}_{\tt in})$ and zero for all other nodes. Given ${\mathcal{G}}({\bm{x}}_{\tt in})$ and the previous (t1)(t-1) steps $V_{1:t-1}=v_{1:t-1}$, the transition probability for the next step is not based on all previous nodes but depends on certain nodes, i.e.
[…]
and zero probability for $\forall v\notin{\mathcal{V}}(v_{1:t-1}\mid{\bm{x}}_{\tt in})$. The sampling process is absorbing upon reaching any node $v\in{\mathcal{V}}_{\tt out}({\bm{x}}_{\tt in})$, indicating that a final answer has been obtained. For non-thinking LLMs, the model directly outputs such types of CoT for the given problem. For thinking LLMs, e.g. DeepSeek-R1 (DeepSeek Team, 2025), the thinking process can be viewed as an exploration of the task-specific DAG with self-correction or backtracking, but its final output shown to the users (excluding thinking tokens) is still consistent with our transition rule.
Applied to our logarithmic count problem, Eq.˜2.1 enforces valid transitions over the nodes. For instance, after collecting $\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$, the next admissible node should be $v_{6}$; while nodes $v_{7}$ through $v_{11}$ remain inaccessible until $v_{6}$ has been visited. We use LLaMA-3.1-8B-Instruct (Grattafiori et al., 2024) to generate four CoT trajectories, where each trajectory consists of its own steps leading to correct/incorrect answers. The nodes and edges shown in Fig.˜1 are performed by the authors, but can also be carried out by LLMs through appropriate prompting (see Section˜4). Accordingly, the LLM’s final CoT output can be split into three classes:
  • Perfect reasoning $(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7},v_{8},{\color[rgb]{0,0.8,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.8,0}v_{10}})$: The trajectory only includes the correct sink node and its ancestors. We formally define this in the next section.
  • Imperfect reasoning [^3] The LLM may reason imperfectly during its thinking process, e.g., dead-ends, self-correction, but is expected to output only the finalized, perfect, correct reasoning results to the user. The reasoning evaluation in this paper is based on the entire final output (while PASS@$k$ just considers the final answer)., e.g., $(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6},v_{7},\boxed{v_{9}},v_{8},{\color[rgb]{0,0.8,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.8,0}v_{10}})$: The trajectory still reaches the correct answer but also includes the irrelevant node $v_{9}$, which is not an ancestor of the correct node. Such case may occur when the LLM explores multiple directions and eventually arrives at the correct answer either by chance or through subsequent derivation. We give an example from AIME 2025 (Art of Problem Solving, 2025a, b) for this case in Appendix˜B where two solution paths are mixed.
  • Wrong reasoning, e.g., $(v_{1},v_{2},v_{3},v_{6},v_{7},v_{9},{\color[rgb]{0.8,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.8,0,0}v_{11}})$: The final answer is incorrect.
Comparison with previous work: Our framework integrates an instance-specific DAG with a rule-based stochastic process, directly addressing key limitations in prior work. We do not assume a fixed “universal" graph with deterministic matching Dziri et al. (2023). Instead, our DAG is logical, acyclic, and features absorbing sink nodes, preserving the directional, goal-oriented nature of mathematical problem solving while allowing for long-range dependencies, in contrast to reversible Markov chain models (Kim et al., 2025). Furthermore, unlike the tree abstractions in Shalev-Shwartz & Shashua (2025), our DAG captures shared sub-derivations and the true dependency structure rather than just linear solution paths, thereby supporting multiple valid routes to a solution.
3 Formal Definition of Mathematical Reasoning Ability
Based on our DAG framework, we now present a formal definition of mathematical reasoning ability. Given an input prompt ${\bm{x}}_{\tt in}$, we independently draw $N$ CoT trajectories $\{{\bm{v}}^{(i)}\}_{i=1}^{N}$ under the proposed sampling mechanism in Eq.˜2.1. For each trajectory ${\bm{v}}^{(i)}$, we construct a trajectory-specific DAG:
${\mathcal{G}}_{\text{gen}}^{(i)}({\bm{x}}_{\tt in})=\bigl({\mathcal{V}}_{\text{gen}}^{(i)}({\bm{x}}_{\tt in}),{\mathcal{E}}_{\text{gen}}^{(i)}({\bm{x}}_{\tt in})\bigr),\quad 1\leq i\leq N\,$,
where the subscript gen indicates that the object (DAG, node, or edge) is extracted from the generated CoT trajectory. Here, ${\mathcal{V}}_{\text{gen}}$ corresponds to the enumerated steps in the trajectory, and ${\mathcal{E}}_{\text{gen}}$ contains edges explicitly defined by the parents of each step. Each trajectory-specific DAG is a sub-DAG of ${\mathcal{G}}({\bm{x}}_{\tt in})$, and the reasoning ability of each trajectory can be evaluated using a new metric, termed logical closeness, and the concept of perfect reasoning, introduced in our framework.
Definition 3.1 (Logical closeness and perfect reasoning).
Under Assumption˜2.1, consider an input prompt ${\bm{x}}_{\tt in}$ and the DAG ${\mathcal{G}}_{\text{gen}}({\bm{x}}_{\tt in})$. For each node $v\in{\mathcal{G}}_{\text{gen}}({\bm{x}}_{\tt in})$, define its out-degree as
${\tt deg}\bigl(v\mid{\mathcal{G}}_{\text{gen}}({\bm{x}}_{\tt in})\bigr):=\Bigl|\bigl\{u\in{\mathcal{G}}_{\text{gen}}({\bm{x}}_{\tt in})\mid(v\rightarrow u)\in{\mathcal{E}}_{\text{gen}}({\bm{x}}_{\tt in})\bigr\}\Bigr|.$
We say that ${\mathcal{G}}_{\text{gen}}({\bm{x}}_{\tt in})$ is logically closed if
${\tt deg}\bigl(v\mid{\mathcal{G}}_{\text{gen}}({\bm{x}}_{\tt in})\bigr)\geq 1,\quad\forall\,v\in{\mathcal{V}}_{\text{gen}}({\bm{x}}_{\tt in}),$
i.e., only the final nodes have no outgoing edges. Furthermore, if the sink node corresponds to the correct answer, we call the associated CoT trajectory a case of perfect reasoning.
Any topological ordering of the ancestor nodes that terminates at the correct sink node is perfect reasoning. Compared with evaluating reasoning based solely on the correctness of the final answer, incorporating logical closeness allows us to assess whether an LLM engages in genuine logical inference rather than merely searching among possible solutions. Based on the definition of logical closeness, we now formally define the mathematical reasoning ability of LLMs as follows.
Definition 3.2 (Mathematical reasoning ability).
Under Assumption˜2.1, let an LLM be given a prompt ${\bm{X}}_{\tt in}\in{\mathcal{P}}$, sampled from an underlying distribution $\mathcal{D}$} over mathematical problem prompts. We define two indicator functions for a trajectory-specific DAG ${\mathcal{G}}_{\text{gen}}({\bm{X}}_{\tt in})$:
$\delta_{\tt close}\bigl({\mathcal{G}}_{\text{gen}}({\bm{X}}_{\tt in})\bigr):=\begin{cases}1,&\text{if ${\mathcal{G}}_{\text{gen}}({\bm{X}}_{\tt in})$ is logically closed},\\ 0,&\text{otherwise},\end{cases}$
$\delta_{\tt final}\bigl({\mathcal{G}}_{\text{gen}}({\bm{X}}_{\tt in})\bigr):=\begin{cases}1,&\text{if the sink node of ${\mathcal{G}}_{\text{gen}}({\bm{X}}_{\tt in})$ is correct},\\ 0,&\text{otherwise}.\end{cases}$
Then, the Perfect Reasoning Rate (PRR) of an LLM w.r.t. a given prompt ${\bm{X}}_{\tt in}$ is defined as
$\operatorname{PRR}({\bm{X}}_{\tt in}):=\mathbb{E}_{{\mathbb{P}}}\Big[\delta_{\tt close}\bigl({\mathcal{G}}_{\text{gen}}({\bm{X}}_{\tt in})\bigr)\times\delta_{\tt final}\bigl({\mathcal{G}}_{\text{gen}}({\bm{X}}_{\tt in})\bigr)\Big].$
The overall mathematical reasoning ability of an LLM over the distribution $\mathcal{D}$} is then measured as
$\mathcal{R}:=\mathbb{E}_{{\bm{X}}_{\tt in}\sim{\mathcal{D}}}\big[\operatorname{PRR}({\bm{X}}_{\tt in})\big]=\mathbb{E}_{{\mathbb{P}},\,{\bm{X}}_{\tt in}\sim{\mathcal{D}}}\Big[\delta_{\tt close}\bigl({\mathcal{G}}_{\text{gen}}({\bm{X}}_{\tt in})\bigr)\times\delta_{\tt final}\bigl({\mathcal{G}}_{\text{gen}}({\bm{X}}_{\tt in})\bigr)\Big].$
AUC socres: By relaxing $\delta_{\tt close}$ to permit a certain proportion of nodes that do not satisfy logical closeness, we obtain the corresponding AUC scores (with proportion of logic closeness from 0% to 100%), which serves as a comprehensive measure of mathematical reasoning performance.
This metric can be regarded as a combination of the final-answer correctness reflected by PASS@1 and logical closeness. To illustrate Definition˜3.2, consider a toy DAG example in Fig.˜2, consisting of two linear chains of length $L$ emanating from a common source node. We denote the correct sink node as L and the incorrect sink node as $\color[rgb]{0.7,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.7,0,0}\textsf{L}^{\prime}$}. At each step, the transition distribution $\mathbb{P}$} is uniform over all available nodes according to Eq.˜2.1, i.e.,
$\forall\,v\in{\mathcal{V}}(v_{1:t-1}\mid{\bm{x}}_{\tt in}),\quad{\mathbb{P}}(V_{t}=v\mid V_{1:t-1}=v_{1:t-1},{\bm{X}}_{\tt in}={\bm{x}}_{\tt in})=\frac{1}{2}.$
To remain logically closed, a trajectory must stay on the same chain across the remaining L1L-1 transitions, each occurring with probability $1/2$. Hence, we have $\operatorname{PRR}=\left(\frac{1}{2}\right)^{L-1}.$ This illustrates that PRR decays exponentially with depth: although the final-answer accuracy ($1/2$) may appear stable, logically closed trajectories become increasingly rare. Furthermore, Appendix˜B presents an example from AIME 2025 (Art of Problem Solving, 2025a, b) where the final answer is correct, but logic-closeness fails due to mixing two solution paths. Consequently, lacking logic closeness risks producing answers correct only by chance, obscuring flawed reasoning, reducing reliability, and undermining interpretability.
Figure 2: Toy DAG.
In practice, given a dataset with $M$ problems and $N$ independent CoT trajectories per problem, one can approximate $\operatorname{PRR}\left({\bm{x}}_{\tt in}\right)$ and $\mathcal{R}$ by
$\widehat{\operatorname{PRR}}\left({\bm{x}}_{\tt in}\right):=\frac{1}{N}\sum_{i=1}^{N}\delta^{(i)}_{\tt close}\left({\bm{x}}_{\tt in}\right)\times\delta^{(i)}_{\tt final}\left({\bm{x}}_{\tt in}\right),\quad\widehat{\mathcal{R}}:=\frac{1}{M}\sum_{j=1}^{M}\widehat{\operatorname{PRR}}\left({\bm{x}}_{\tt in}^{(j)}\right).$
By Hoeffding’s inequality, taking $M=\Omega(1/\varepsilon^{2})$ is sufficient for $\big|\widehat{\mathcal{R}}-\mathcal{R}\big|\leq\varepsilon$ with high probability.
Comparison with generalization in supervised learning: Analogous to underfitting and overfitting in supervised learning, we can define under-reasoning-where the DAG omits necessary intermediate steps-and over-reasoning-where the DAG is logically sound but contains redundant or irrelevant steps. In both cases, the estimated reasoning ability $\widehat{\mathcal{R}}$ is low, and perfect reasoning can be regarded as the “sweet spot” in this reasoning-generalization analogy.
Importantly, $\mathcal{R}$ offers a richer error taxonomy, distinguishing structurally illegal paths, legal-but-wrong paths, and “imperfectly correct” trajectories, whereas standard generalization theory typically treats all errors uniformly. Regularization strategies, such as the minimum description length principle (Grünwald, 2007), could potentially mitigate over-reasoning by favoring concise proofs that conform to a proof grammar or template. We leave exploration of this direction to future work.
4 DAG-MATH Formatted CoT and Benchmark
In practice, standard CoT trajectories generated by LLMs are unstructured, autoregressive sequences of tokens. Within such free-form text, the logical steps (nodes) and their dependencies (edges) are often entangled, which complicates the evaluation of our step-level reasoning framework. To address this issue, we introduce a structured CoT format via prompting, described in Section˜4.1, which we term the DAG-MATH format. This format facilitates the construction of the corresponding DAG, enabling the creation of a DAG benchmark, presented in Section˜4.2.
4.1 DAG-MATH Formatted CoT
As a structured, step-by-step format, DAG-MATH explicitly specifies each reasoning step in forward generation order: Edge $\rightarrow$ Parent(s) $\rightarrow$ Node. This ordering is designed for evaluation-oriented sampling: first, declare the logical link to prior knowledge (Edge); next, cite the necessary antecedents (Parents); and finally, assert the derived conclusion (Node). The DAG-MATH format can be produced by LLMs through prompt engineering. For illustration, below is one step from the DAG-MATH formatted CoT for the logarithmic count problem (all steps in DAG-MATH format are presented in Appendix˜C).
Step 4 Edge: Since the left-hand side of the equation in Step 1 contains $\log(x-1)$, the domain restriction for a logarithm requires x1>0x-1>0, i.e., $x>1$. Parents: Step 1.         Node: $\log(x-1)$ requires $x>1$.
Step IDs also serve as node identifiers, which allows for a straightforward evaluation of DAG closeness. Based on the DAG-MATH format, a gold-standard CoT trajectory is defined by three criteria: (1) it adheres to the DAG-MATH format; (2) its corresponding DAG is logically closed; and (3) the final answer is correct.
4.2 Gold-Standard DAG-MATH Benchmark
We prompt LLMs to generate CoT trajectories in the DAG-MATH format for existing mathematical datasets, such as Omni-MATH (Gao et al., 2024), and construct the corresponding DAGs. By verifying both logical closeness (Definition˜3.1) and the correctness of the final answers, we compile a benchmark consisting of 2,894 gold-standard DAGs. The primary purpose of this benchmark is to characterize the statistical properties of these gold-standard DAGs across different problem difficulty levels, providing valuable insights for evaluating and enhancing LLM mathematical reasoning.
The benchmark comprises problems from Omni-MATH (Gao et al., 2024), which are categorized into difficulty levels ranging from 1 (easiest) to 10 (hardest). To ensure high solvability by LLMs, we only consider problems with difficulty levels below 6. For generating DAG-MATH formatted CoTs, we employ GPT-o4-mini and Qwen3-235B-A22B-Thinking-2507, both recognized as leading models in mathematical tasks (LMArena, 2025). Gold-standard CoTs are constructed using a three-stage prompting strategy (see Appendix˜D) in reverse order (Node $\rightarrow$ Parents $\rightarrow$ Edge). This approach fixes the node set first, making verification easier with SymPy or LLM-as-Judge and minimizing error propagation, thereby ensuring high-quality trajectories.
We consider five representative graph-level statistics: (1) the total number of nodes (#Nodes); (2) the total number of edges (#Edges); (3) graph density, defined as the ratio of #Edges to the maximum possible number of edges in an acyclic graph, i.e., $\frac{2\#\text{Edges}}{\#\text{Nodes}(\#\text{Nodes}-1)}$; (4) the maximum in-degree, denoted ${\operatorname{d}}^{\max}_{\tt in}$; and (5) the maximum out-degree, denoted ${\operatorname{d}}^{\max}_{\tt out}$. Fig.˜3 shows the distributions of these five statistics across problem difficulty levels. Our key observations as problem difficulty increases from 0 to 6 are as follows:
  • More nodes and edges with heavier tails: The distributions of #Nodes and #Edges shift noticeably to the right and develop heavier tails, indicating that harder problems produce larger graphs, while simpler problems yield much smaller ones.
  • Sparser structure: The graph becomes sparse when problem difficulty increases. Harder reasoning produces broader, less connected structures, reflecting modular sub-reasoning where semi-independent chains (e.g., sub-tasks or lemmas) are later combined.
  • Logic complexity reflected in maximum out-degree: As difficulty increases, the distributions of maximum in-degree and out-degree shift rightward with heavier tails. Maximum in-degree grows slowly, suggesting most steps rely on few inputs, whereas maximum out-degree rises more sharply, indicating that certain key steps support multiple inferences. This implies that logical complexity scales primarily through branching rather than aggregation. The average in- and out-degree remains around 1.3 across difficulty levels, as most nodes have small degrees while a few pivotal steps exhibit large connectivity.
Accordingly, as problems become harder, their DAGs grow larger and sparser, with complexity arising primarily from branching into modular sub-reasoning chains. This underscores the importance of LLMs being able to decompose problems into sub-tasks, track longer dependencies, and recombine intermediate results to solve challenging problems effectively.
Figure 3: Empirical distributions of DAG statistics across problem difficulty groups, where group $k$ corresponds to problems with difficulty in (k1,k](k-1,k]. Shown are the distributions of #Nodes, #Edges, graph density $\left(\text{i.e., }\frac{2\#\text{Edges}}{\#\text{Nodes}(\#\text{Nodes}-1)}\right)$, maximum in-degree, and maximum out-degree.
5 Evaluation of Mathematical Reasoning Ability
To evaluate mathematical reasoning performance, we employ few-shot prompting (using demonstrations from the benchmark in Section˜4.2; see Appendix˜E for details) to guide LLMs in generating DAG-MATH formatted CoTs/DAGs for test problems, without providing the final solutions.
Models and datasets: We evaluate five LLMs: Gemini-2.5-Flash (Gemini-2.5-F), Gemini-2.5-Flash-Lite (Gemini-2.5-F-L), GPT-4.1, GPT-4.1-mini (GPT-4.1-M), and Qwen3-30B-A3B-Instruct-2507 (Qwen3-30B). These models demonstrate strong mathematical performance even without long thinking (White et al., 2025), also are more efficient and economical than other thinking-focused models due to lower token usage. We evaluate these models on three recently adopted datasets for high-difficulty, answer-based problems: AIME 2025 (Art of Problem Solving, 2025a, b), BRUMO 2025 (BRUMO, 2025), and HMMT 2025 (HMMT, 2025).
Experimental Settings: For each model, we use a 4-shot prompting strategy to generate 32 DAG-MATH formatted CoT trajectories per problem across all datasets. Next, we extract the corresponding DAGs and evaluate the five graph-level statistics introduced in Section˜4.2, as well as model performance metrics (e.g., PASS@1 and $\widehat{\mathcal{R}}$). The sampled DAGs are then partitioned into four classes: All (no filtering), Incorrect (ending at an incorrect sink), Correct (ending at the correct sink), and Perfect (logically closed and ending at the correct sink).
Fig.˜4 reports AUC scores across three datasets, with PASS@1 as the starting point and $\widehat{\mathcal{R}}$ as the end point (see more details in Table˜2 in Appendix˜F), as the logic closeness rate increases. Besides, we also report averaged graph-level statistics for AIME 2025 in Table˜1, with additional results for BRUMO 2025 and HMMT 2025 in Appendix˜F. We have the following observations:
Figure 4: The AUC curves of averaged accuracy under the logical closeness rate over three datasets for five selected LLMs.
Table 1: Averaged graph-level statistics of sampled DAGs across selected LLMs on AIME 2025.
Model Class #nodes #edges density ${\operatorname{d}}^{\max}_{\tt in}$ ${\operatorname{d}}^{\max}_{\tt out}$
Gemini-2.5-F All 32.8 48.9 11.2% 4.3 7.0
Incorrect 35.6 53.3 10.6% 4.6 8.6
Correct 30.2 45.1 11.6% 4.1 5.5
Perfect 23.3 30.8 13.0% 3.3 3.6
Gemini-2.5-F-L All 33.0 54.0 13.4% 3.6 9.7
Incorrect 40.5 68.6 11.9% 3.9 12.8
Correct 21.5 31.6 15.7% 3.2 4.8
Perfect 16.1 21.4 18.4% 3.0 3.2
GPT-4.1 All 17.8 21.4 16.2% 2.6 3.0
Incorrect 18.4 22.2 15.6% 2.6 3.1
Correct 15.9 19.3 17.5% 2.7 2.9
Perfect 14.1 16.8 19.0% 2.5 2.4
GPT-4.1-M All 22.8 31.3 14.2% 3.2 4.0
Incorrect 25.0 34.7 13.3% 3.3 4.3
Correct 17.9 23.6 16.6% 3.0 3.4
Perfect 16.5 21.5 17.4% 3.0 3.1
Qwen3-30B All 21.4 31.1 16.0% 3.3 4.9
Incorrect 23.4 34.3 14.9% 3.4 5.5
Correct 18.9 27.2 17.2% 3.1 4.1
Perfect 14.7 19.6 20.2% 2.9 3.0
  • Search improves raw accuracy while perfect reasoning ability remains similar. All models exhibit a noticeable drop from PASS@1 to $\widehat{\mathcal{R}}$; while PASS@1 varies widely across models, $\widehat{\mathcal{R}}$ remains relatively stable (i.e., the end point is almost the same). This suggests that additional exploration or search can inflate raw accuracy, while the models’ inherent perfect reasoning ability is broadly comparable. The AUC scores indicate that models with correct answers achieve at most 80% logic-closed nodes, but accuracy degrades markedly under stricter criteria. This suggests that outputs, while correct, are superficially consistent at some point, which aligns with users’ impressions when using these LLMs.
  • Graph structure reflects problem difficulty and reasoning quality. Harder problems induce larger, sparser, and more branchy DAGs (see Section˜4.2). Within each model, Perfect reasoning trajectories correspond to the smallest, densest graphs, reflecting concentrated reasoning on solvable items. Correct graphs are slightly larger and sparser, suggesting the inclusion of useful exploratory steps. In contrast, Incorrect graphs exhibit strong branching (with $\widehat{d}^{\max}_{\tt out}$ growing faster than $\widehat{d}^{\max}_{\tt in}$), indicating that failure often arises from speculative expansions rather than from aggregating insufficient inputs. For example, Gemini’s Correct cohort exhibits larger but sparser graphs with slightly higher branching when compared to the GPT family, indicating that effective exploration and task planning can increase the likelihood of reaching correct answers without fully closed, deeper reasoning.
  • Identifying the “difficulty boundary”. Within each model, the Incorrect cohorts resemble “harder-than-ability” graphs (see Section˜4.2): larger, sparser, heavy-tailed, and with notably higher $\widehat{d}^{\max}_{\tt out}$. In contrast, the Perfect cohorts converge to smaller, relatively dense DAGs with low branching. This indicates that each model’s effective difficulty ceiling corresponds to the regime where it can maintain compact, low-branching DAGs; beyond this point—when branching explodes and density drops—accuracy sharply declines.
6 Conclusion
This paper proposes a novel DAG-MATH framework for modeling and evaluating mathematical reasoning, introducing the concepts of logical closeness and perfect reasoning over DAGs. We demonstrate how DAG graph statistics vary with problem difficulty and how models’ perfect-reasoning ability and AUC curve behaves across these tasks. The framework provides an accessible mathematical formalization of reasoning and memorization in LLMs, paving the way for future work on reasoning guarantees, analogous to generalization guarantees in supervised learning.
Acknowledgment
Y. Z. was supported by Warwick Chancellor’s International Scholarship. JDL acknowledges support of Open Philanthropy, NSF IIS 2107304, NSF CCF 2212262, ONR Young Investigator Award, NSF CAREER Award 2144994, and NSF CCF 2019844. F. L. was supported by Royal Society KTP R1 241011 Kan Tong Po Visiting Fellowships and Warwick-SJTU seed fund. We thank Zulip [^4] https://zulip.com/ for the project organization tool and Sulis [^5] https://warwick.ac.uk/research/rtp/sc/sulis/ for GPU computation resources.
Appendix A Illustration and Definition of CoT Step
In this section, we first discuss execution correctness as a measure of LLM reasoning performance and emphasize that our framework addresses the key challenge of tracking logical dependencies based on a formal mathematical definition of CoT.
A.1 Remark on Execution Correctness
We note that accurate calculation and symbolic execution remain essential for evaluating an LLM’s mathematical reasoning. In practice, LLMs may make stepwise errors, such as computational slips or misreading problem statements. Our framework, however, assumes that each step is correct. This simplification is justified because step-level errors can be automatically detected using Process Reward Models (Lightman et al., 2023; Wang et al., 2024; Zhang et al., 2025) or SymPy validation.
Focusing on correct steps allows us to capture the trajectory of model capabilities: with recent advances (Kavukcuoglu, 2025; OpenAI, 2025; Yang et al., 2025; DeepSeek Team, 2025), the main bottleneck in complex mathematical reasoning lies less in local step fidelity and more in the higher-level logical structure. Our framework addresses this challenge by: (1) perceiving the full logical structure of a problem, and (2) navigating it to construct a coherent solution path. By abstracting away from local step errors, we can isolate and analyze the structural core of CoT reasoning.
A.2 Mathematical Definition of Steps in CoT
We now formalize the definition of a single CoT step. Let $\mathcal{V}$} denote the semantic domain of mathematical objects in semantic normal form (SNF). A raw CoT step is a sequence of tokens, i.e., a string ${\bm{c}}\in{\mathcal{P}}$. We introduce a canonicalization map
$\kappa:{\mathcal{P}}\to{\mathcal{V}}$
that maps any textual span to its SNF representation:
${\bm{c}}\;\xrightarrow{\;\kappa\;}\;\kappa({\bm{c}})\in{\mathcal{V}}\,.$
The canonicalization map satisfies the following properties:
  • Idempotent: $\kappa(\kappa({\bm{c}}))=\kappa({\bm{c}})$.
  • Presentation-invariant: It does not perform substantive algebraic manipulations (e.g., expanding or factoring), which are treated as separate reasoning steps.
Canonicalization removes superficial variations such as synonyms, spacing differences, commutativity, and $\alpha$-equivalent variable names, ensuring semantic consistency.
Intuitively, a CoT step can be mapped to a normalized SymPy object that captures its underlying mathematical semantics. A concrete example of this canonicalization is provided below.
  • Input: Raw CoT step as text tokens (e.g., “The value of our target is x+yx+y.” or “The target value is (y+x)(y+x).”).
  • Canonicalization ($\kappa$): Maps the input to a normalized form by removing superficial variations such as synonyms, spacing, commutativity, and $\alpha$-equivalent variable names.
  • Output: A semantic object (e.g., SymPy object Add(x,y)) that consistently encodes the meaning.
Accordingly, let $\mathcal{F}$ denote a signature of primitive inference rules/operations. Each $f\in\mathcal{F}$ is associated with an arity $\operatorname{ar}(f)\in\mathbb{N}$ (i.e., the number of inputs the rule or operation takes) and a (partial) semantic operator $\llbracket f\rrbracket:{\mathcal{V}}^{\operatorname{ar}(f)}\rightharpoonup{\mathcal{V}}$. Intuitively, $\mathcal{F}$ represents the atomic reasoning steps allowed at the CoT level, such as a single algebraic operation, one application of a named lemma, or a single substitution. We now formally define a CoT step.
Definition A.1 (Atomic Step of CoT).
Given an input prompt ${\bm{x}}_{\tt in}\in{\mathcal{P}}$, a CoT trajectory of length $\ell$ is a sequence of steps $\mathcal{C}=(\mathbf{c}_{1},\dots,\mathbf{c}_{\ell})$. Suppose $\kappa:{\mathcal{P}}\to{\mathcal{V}}$ is a canonicalization map. Then, a step $\mathbf{c}_{i}$ can be formulated as a triple $(\Gamma_{i},f_{i},v_{i})$ denoted by $\Gamma_{i}\overset{f_{i}}{\longrightarrow}v_{i}$, where:
  1. 1.
    Canonicalization: Each string ${\bm{c}}_{i}\in{\mathcal{P}}$ produced in the CoT trajectory is mapped by $\kappa$ to the corresponding SNF object $v_{i}=\kappa({\bm{c}}_{i})\in{\mathcal{V}}$.
  2. 2.
    Premises: $\Gamma_{i}\subseteq\{v_{1},v_{2},\dots,v_{i-1}\}$ is the finite set of previously established SNF objects (from the prompt or earlier steps) directly used to infer the current step.
  3. 3.
    Primitive operation: $f_{i}\in\mathcal{F}$ and $v_{i}=\llbracket f_{i}\rrbracket(\Gamma_{i})$, i.e., $v_{i}$ is obtained by exactly one application of a primitive operator to the premises.
Accordingly, each step in a CoT can be viewed as the reasoning pattern:
$\text{(Premises used)}+\text{(inference rule applied)}\;\longrightarrow\;\text{(new result)}.$
Next, we provide a concrete algebra example on expanding $(x+y)^{2}$ for better intuition.
Step $c_{i}$: Expand the square
$\Gamma_{i}=\{(x+y)^{2}\},\quad f_{i}=\text{``expand square''},\quad v_{i}=\llbracket f_{i}\rrbracket(\Gamma_{i})=(x+y)(x+y).$
So we have:
$\{(x+y)^{2}\}\;\xrightarrow{\text{expand square}}\;(x+y)(x+y).$
Step $c_{i+1}$: Distribute the product
$\Gamma_{i+1}=\{(x+y)(x+y)\},\quad f_{i+1}=\text{``distributive law''},\quad v_{i+1}=\llbracket f_{i+1}\rrbracket(\Gamma_{i+1})=x^{2}+xy+yx+y^{2}.$
So we have:
$\{(x+y)(x+y)\}\;\xrightarrow{\text{distribute}}\;x^{2}+xy+yx+y^{2}.$
Step $c_{i+2}$: Simplify like terms
$\Gamma_{i+2}=\{x^{2}+xy+yx+y^{2}\}\,$,
$f_{i+2}=\text{``commutativity + combine like terms''}\,$,
$v_{i+2}=\llbracket f_{i+2}\rrbracket(\Gamma_{i+2})=x^{2}+2xy+y^{2}\,.$
So we have:
$\{x^{2}+xy+yx+y^{2}\}\;\xrightarrow{\text{simplify}}\;x^{2}+2xy+y^{2}.$
Combining the above steps, the CoT trajectory is
$(x+y)^{2}\;\xrightarrow{\,\text{expand square}\,}\;(x+y)(x+y)\;\xrightarrow{\,\text{distribute}\,}\;x^{2}+xy+yx+y^{2}\;\xrightarrow{\,\text{simplify}\,}\;x^{2}+2xy+y^{2}.$
This example follows Definition˜A.1 and precisely characterizes a CoT trajectory at the step level. Note that this step-level formalization is not essential for understanding the main text, which primarily focuses on DAG-level reasoning rather than the specifics of individual nodes and edges. Nonetheless, for readers interested in how nodes and edges are defined or how they influence a CoT trajectory, this definition and the accompanying example provide a useful reference.
Appendix B Example of Logical Closeness
There are several reasons why LLMs may generate unclosed nodes even though the final answer is correct:
  • Assertions stemming from an alternative strategy that is not the one leading to the final answer in the trajectory.
  • Qualitative axioms that are implicitly used. When forming edges, the model tends to link parents that provide numerical values from earlier calculations, since quantitative conclusions are easier to cite than qualitative ones.
  • Irrelevant information drawn from the problem statement.
  • Additional commentary based on previous conclusions but not required for the solution.
We aim to analyze specific DAG-MATH formatted CoT trajectory which has the correct final answer but unclosed DAG. To justify the rationale, we take the following geometry problem from AIME 2025 I (Art of Problem Solving, 2025a) as an example.
Area of Heptagon Problem On $\triangle ABC$, points $A,D,E$, and $B$ lie in that order on side $\overline{AB}$ with $AD=4$, $DE=16$, $EB=8$. Points $A,F,G$ and $C$ lie in that order on side $\overline{AC}$ with $AF=13$, $FG=52$, and $GC=26$. Let $M$ be the reflection of $D$ through $F$, and let NN be the reflection of $G$ through $E$. Quadrilateral DEGFDEGF has area $288$. Find the area of heptagon AFNBCEMAFNBCEM.
We have $4$ correct CoT trajectories over $32$ total trajectories generated by Gemini-2.5-Flash. We provide a detailed analysis of one trajectory that contains multiple unclosed patterns and has a moderate graph size. There are two trajectories that exhibit similar characteristics, while the last trajectory consists of $121$ nodes.
[…]
calc
{tikzpicture}
[scale=0.045, line join=round, line cap=round, >=stealth]
[…]
(A) at (100,100);
[…]
[draw=black, fill=gray!20] (N) – (E) – (M) – (F) – cycle; [draw=black, fill=gray!20] (N) – (E) – (C) – (B) – cycle; [draw=black, fill=gray!20] (A) – (F) – (M) – cycle;
[…]
[line width=0.5mm] (A) – (B) – (C) – cycle;
[…]
Figure 5: Visualization of the heptagon problem.
Figure 6: A trajectory-specific DAG for the Area of Heptagon problem, whose DAG-MATH–formatted CoT trajectory is generated by Gemini-2.5-Flash and has the correct final node. Nodes without any children are highlighted in orange.
We plot Fig.˜5 for better illustration and present the first trajectory-specific DAG shown in Fig.˜6. We summarize each node’s conclusion below.
  1. 1.
    States the ordering/collinearity A,D,E,BA\,,D\,,E\,,B on $\overline{AB}$.
  2. 2.
    States $AD=4$.
  3. 3.
    States $DE=16$.
  4. 4.
    States $EB=8$.
  5. 5.
    Computes $AE=AD+DE=20$.
  6. 6.
    Computes $DB=DE+EB=24$.
  7. 7.
    Computes $AB=AD+DE+EB=28$.
  8. 8.
    States the ordering/collinearity AFGCA-F-G-C on $\overline{AC}$ (global setup).
  9. 9.
    States $AF=13$.
  10. 10.
    States $FG=52$.
  11. 11.
    States $GC=26$.
  12. 12.
    Computes $AG=AF+FG=65$.
  13. 13.
    Computes $FC=FG+GC=78$.
  14. 14.
    Computes $AC=AF+FG+GC=91$.
  15. 15.
    Defines $M$ as the reflection of $D$ across $F$ (so $F$ is midpoint of $DM$).
  16. 16.
    Elaborates reflection at $F$; observes $F$ midpoint of $DM\Rightarrow\text{Area}(AFM)=\text{Area}(ADF)$.
  17. 17.
    Defines $N$ as the reflection of $G$ across $E$ (so $E$ is midpoint of $GN$).
  18. 18.
    Notes explicitly that $E$ is the midpoint of $GN$.
  19. 19.
    States $\text{Area}(DEGF)=288$.
  20. 20.
    Sets $\text{Area}(ADF)=\tfrac{AD}{AB}\cdot\tfrac{AF}{AC}\cdot\text{Area}(ABC)=\tfrac{S}{49}$.
  21. 21.
    Sets $\text{Area}(AEG)=\tfrac{AE}{AB}\cdot\tfrac{AG}{AC}\cdot S=\tfrac{25S}{49}$.
  22. 22.
    Relates $\text{Area}(DEGF)=\text{Area}(AEG)-\text{Area}(ADF)=\tfrac{24S}{49}$.
  23. 23.
    Solves $\tfrac{24S}{49}=288\Rightarrow S=\text{Area}(ABC)=588$.
  24. 24.
    Computes $\text{Area}(ADF)=S/49=12$.
  25. 25.
    Uses reflection at $F\Rightarrow\text{Area}(AFM)=\text{Area}(ADF)$.
  26. 26.
    Concludes $\text{Area}(AFM)=12$.
  27. 27.
    Computes $\text{Area}(ABG)=\tfrac{AG}{AC}\cdot S=420$.
  28. 28.
    Computes $\text{Area}(BEG)=\text{Area}(ABG)-\text{Area}(AEG)=420-300=120$.
  29. 29.
    Uses reflection at $E\Rightarrow\text{Area}(BEN)=\text{Area}(BEG)$.
  30. 30.
    Concludes $\text{Area}(BEN)=120$.
  31. 31.
    Chooses coordinates $A=(0,0),B=(28,0),C=(0,42)$; derives $D,E,F,G,M,N$ coordinates.
  32. 32.
    Applies the shoelace formula to the heptagon $AFNBCEM$ and gets area $588$.
  33. 33.
    States the final result: $\boxed{588}$.
We can observe that the DAG has 8 non-closed nodes. We diagnose the reasons of uncloseness for each node:
  • Nodes $1$ & $8$: These two nodes state the global setup on collinearity/ordering, which are directly provided by the problem statement. The later steps implicitly use collinearity to add segment lengths on $\overline{AB}$ and $\overline{AC}$, but the LLM does not recognize that it has used these two nodes in its subsequent reasoning.
  • Nodes $6$ & $13$ & $16$ & $18$: These nodes derive or state extra commentary of their previous step, which are not needed in the subsequent steps.
Then, the message conveyed by Nodes $26$ & $30$ is crucial. In this problem, the area of the heptagon $AFNBCEM$ can be computed via two distinct strategies:
  • Reflection-swap strategy: The core idea is to replace two interior triangles ($\triangle ADF$, $\triangle BEG$) of $\triangle ABC$ with their exterior reflected counterparts ($\triangle AFM$, $\triangle BEN$), showing that the net area change is zero. Consequently, the heptagon’s area is obtained by a straightforward “remove + add" bookkeeping.
  • Shoelace strategy: This coordinate-based, algebraic method requires only listing the vertices $A,F,N,B,C,E,M$ in order and then applying the determinant sums to compute the area.
Nodes $20$ through $30$ derive the areas required for the final “remove + add” computation in the reflection-swap strategy, namely
$\text{Area}(AFNBCEM)$ $=\underbrace{\text{Area}(ABC)}_{\text{{\bf Node} 23}}\;-\;\underbrace{\text{Area}(ADF)}_{\text{{\bf Node} 24}}\;-\;\underbrace{\text{Area}(BEG)}_{\text{{\bf Node} 28}}$
$\quad\;+\;\underbrace{\text{Area}(AFM)}_{\text{{\bf Node} 26}}\;+\;\underbrace{\text{Area}(BEN)}_{\text{{\bf Node} 30}}\,.$
If the model had continued with this strategy, Node 31 would correspond to the above equation, with Nodes $26$ & $30$ as its parents. However, the model instead switches to the shoelace strategy at Node 31 and successfully obtains the correct answer, leaving Nodes $26$ & $30$ unclosed.
This provides evidence that the model generates elements of an alternative strategy that remain unused in the current trajectory and switches strategies during the generation process.
Next, the second and third trajectories exhibit node structures similar to the first. They also contain nodes such as Nodes $1$ & $8$ in the first trajectory, which are not recognized as being used. However, unlike the first trajectory, they rely solely on the reflection-swap strategy to obtain the final answer without switching strategy.
The final trajectory consists of $121$ nodes in total. We provide a comprehensive review of its reasoning process: it begins by copying the givens, constructing segment sums, and recording the reflection, analogous to Nodes 1–19 in the first trajectory. It then attempts a parametric area strategy via trigonometric parametrization but halts after approximately $20$ steps. Subsequently, it searches over many polygon decompositions of heptagon $AFNBCEM$, repeatedly proposing and discarding formulas—clear evidence of exploratory search—until it identifies the structural invariant $AD:DE:EB=AF:FG:GC=1:4:2$, from which it correctly infers $DF\parallel EG\parallel BC$. At this stage, the strategy shifts to a similarity/area-ratio approach for triangles sharing angle $A$ and successfully derives the area of $\triangle ABC$. Finally, the strategy switches once more to the reflection-swap method, yielding the correct answer.
Appendix C Example DAG-MATH Formatted CoT
Example DAG-MATH Formatted CoT for the logarithmic count problem Step 1 Edge: Restate the target equation from the problem statement, $2\log(x-1)=\log k$, which is the central equation to solve for $x$ in terms of $k$. Parents: Null Node: $2\log(x-1)=\log k$. $\vdots$ Step 4 Edge: Since the left-hand side of the equation in Step 1 contains $\log(x-1)$, the domain restriction for a logarithm requires x1>0x-1>0, i.e., $x>1$. Parents: Step 1 Node: $\log(x-1)$ requires $x>1$. $\vdots$ Step 8 Edge: Using Step 7, the candidate solutions are $1\pm\sqrt{k}$; Step 4 requires $x>1$, and Step 5 ensures $k>0$, so $\sqrt{k}>0$, making $1-\sqrt{k}<1$ invalid. Therefore, $1+\sqrt{k}$ is the only admissible solution. Parents: Step 4,5,7 Node: $1+\sqrt{k}$ is the only admissible solution. Step 9 Edge: From Step 8, each positive integer $k$ yields exactly one solution $x=1+\sqrt{k}$. Step 5 requires $k>0$, and Step 2 restricts $k$ to integers in $[1,300]$. Therefore, there are 300 valid $k$ values, satisfying Step 3’s requirement of exactly one solution. Parents: Step 2,3,5,8 Sink Node: There are $\boxed{300}$ valid values for $k$.
Appendix D Benchmark Construction
In this section, we detail our three-stage strategy for benchmark construction.
Stage 1: We prompt GPT-o4-mini to generate only the Node set, one step at a time, using the instructions in Section˜D.1. To enhance correctness, we adopt a supervised setup that provides both the problem and its correct solution. Because standard CoT can skip arithmetic or combine multiple results in a single step, we require that each Node consist of exactly one sentence containing a single mathematical or logical assertion (i.e., one primitive action per step) to normalize granularity. The complete Node set is then validated using SymPy and LLM-as-Judge; if the final answer or any intermediate assertion is incorrect, the Nodes are resampled and re-validated.
Stage 2: Given the verified Node set from Stage 1, we prompt GPT-o4-mini (per Section˜D.2) to assign, for each Node, a minimal set of direct Parents sufficient to derive it via a primitive operation. We enforce acyclicity and well-typed arity constraints, then assemble the full DAG. The resulting DAG is checked for logical closeness relative to the sink node; if checks fail, dependencies are resampled. Simultaneously, irrelevant Nodes flagged in Stage 1 are pruned so that non-contributing leaves do not persist in the gold graph.
Stage 3: Conditioned on the generated [Parent(s), Node] pairs, we prompt Qwen3-235B-A22B-Thinking-2507666We choose Qwen3 as it provides a clearer description of the inference from Parents to Nodes than GPT-o4. (per Section˜D.3) to generate the Edge content that justifies how each Node is inferred from its Parents. The justification must introduce no new facts beyond the problem statement and the cited Parents. After this step, the triplets are merged into a gold-standard DAG-MATH formatted CoT in forward order.
D.1 Prompt for Stage 1
Stage 1 - Instructions for Refined Step-by-Step Answer Task:
You are an expert in mathematical logic and reasoning. Your job is to take rough multi-step math solutions and rewrite them in a detailed, structured, and logical step-by-step format. Follow these guidelines:
[…]
D.2 Prompt for Stage 2
Stage 2 - Instructions for Step Dependency Analysis Role and Objective
You are an expert in mathematical logic and stepwise reasoning. Your primary role is to analyze the detailed solution steps for a mathematical problem and annotate each step with its minimal set of direct dependencies.
Instructions 1. Input Structure: Each problem appears as a JSON object with fields: problem_text (string): Problem description. final_answer (string): The solution. steps (array): Each is an object containing: step_id (integer): Unique step identifier. text (string): Reasoning or mathematical operation. 2. Dependency Annotation: Add a direct_dependent_steps field for every step. For each step: If stated directly from the problem statement, assign null. Otherwise, list the minimal, directly required prior step_id values in ascending order, e.g., [2, 3]. Dependency Rules: 1. Every dependency ID in direct_dependent_steps MUST exist in the original set of step_id values of the input. 2. Every listed dependency must be a prior step—its step_id must be strictly less than the current step (i.e., no self- or future-dependency is allowed). 3. Self-Validation for Step Closure: After assigning dependencies, check for unclosed intermediate steps: An unclosed step is any non-final step not referenced in
direct_dependent_steps by any subsequent step.
If any exist, refine dependencies until all intermediate steps are "closed" (each is used at least once by a later step).
4. Post-action Validation: After annotating dependencies and closing all steps, validate that each intermediate step is referenced at least once by a subsequent step before finalizing the output. If any issues are found, self-correct and repeat the closure process. 5. Structured Output and Consistency: The number of steps in the structured output MUST match exactly the number of steps in the original input.
Application Process When presented with a new problem in valid JSON: 1. Iterate through steps in order of step_id. 2. For each step, determine if it relies on previous steps or the problem statement and annotate direct_dependent_steps as specified. 3. Validate dependencies (all dependency IDs exist; all point to a prior step). 4. Check for unclosed steps and adjust as necessary. 5. Validate closure and present the final modified problem. Error Handling If the input is malformed (e.g., required fields missing, step_id missing, or step_id values not strictly ascending): Return only a JSON object with an error field and a concise message. For example: "error": "Input data malformed: missing step_id in step 3." Output Format Output the structured response with all steps mirroring the original order and count, each annotated with direct_dependent_steps. If there is an error, output only the error object as described. {JSON text of refined step-by-step answer with problem statement from Stage 1}
D.3 Prompt for Stage 3
Stage 3 - Instructions for Constructing Edge Inference Role and Objective You serve as a mathematics solution explainer. For each step in a solved mathematics problem (provided as structured JSON), generate a detailed and explanatory justification paragraph (edge string) clarifying the correctness and logical progression. Instructions For each step in the provided "steps" array within the problem JSON, write a edge string that satisfies the rules below. If direct_dependent_steps is not null, explicitly justify how
direct_dependent_steps support the current one, citing their IDs. If null, note that the step is given by the problem statement or background knowledge (definition, theorem, lemma, fact, or general knowledge not originally in the problem statement).
Clearly explain the mathematical or logical principle, operation, or procedure used (e.g., counting rule, algebraic manipulation, definition, theorem, arithmetic operation) applied in the step. Strict Rule: Do not omit any referenced dependency, the edge MUST include every step in direct_dependent_steps. Clearly illustrate why we need to do this step and how we arrive at this step (acts like planning). For numeric calculations, perform the arithmetic clearly and include a brief sanity check as appropriate. Use neutral, present-tense, explanatory sentences with active voice. Ensure each justification is self-contained, needing only the current step and referenced prior steps to be understandable.
Context Input: JSON object representing a solved math problem with an array of steps, each with step_id, text, and direct_dependent_steps. Output: Return only the edge for each step, ordered to match the original steps array. Do NOT include internal model reasoning or any execution commentary in the output. Demonstration Example Here is a demonstration of the style and format of the edge field. You need to follow this style and format. “We plan to exclude numbers divisible by 2, 3, or 5. To do that systematically we first express the count of multiples of each relevant divisor in the domain. The number of multiples of $k$ up to $n$ is $\left\lfloor\frac{n}{k}\right\rfloor$. Applying that with $n=999$ (from Step 1) and $k=2$ gives $\lfloor 999/2\rfloor$. Writing it as a floor expression is precise and handles the non-divisible endpoint gracefully." “Since numbers divisible by 2, 3, and 5 are overlapping, we need to subtract the count of numbers divisible by pairwise combinations of 2, 3, and 5 then add back the count of numbers divisible by all three, i.e., inclusion-exclusion. The least common multiple of 2 and 3 is 6, so count multiples of 6. Recall we have 999 integers in the domain from Step 1, count multiples of 6 up to 999 via $\lfloor 999/6\rfloor$. This uses the same floor-division approach but applied to the least common multiple of the pair." “We convert the expression in Step 8 into a concrete number to use in later arithmetic. Compute $999/6=166.5$; taking the floor yields $166$. Quick cross-check: $166\times 6=996$, so the last multiple is exactly 996." “We now have all the building blocks: counts of singles, pairwise intersections, and the triple intersection. The immediate plan is to apply the inclusion-exclusion formula for three sets to compute the size of the union $A\cup B\cup C$ where $A=$ multiples of 2, $B=$ multiples of 3, $C=$ multiples of 5. Inclusion-exclusion avoids overcounting and is the rigorous combinatorial tool for unions of overlapping sets. The three-set inclusion-exclusion identity is $|A\cup B\cup C|=|A|+|B|+|C|-(|A\cap B|+|A\cap C|+|B\cap C|)+|A\cap B\cap C|$. Substitute the numerical values computed from Steps 3, 5, 7, 9, 11, 13, 15,: 499+333+199(166+99+66)+33499+333+199-(166+99+66)+33. Writing it explicitly as 499+333+1991669966+33499+333+199-166-99-66+33 lays out the arithmetic to be performed next." “Cite the standard definition: 1 has exactly one positive divisor (1 itself), so it is neither prime (requires two distinct divisors) nor composite (requires more than two). In partitioning the 266 numbers not divisible by 2,3,5 into primes and composites, we must also handle the special case 1, which is counted among the 266 but is neither prime nor composite; failing to account for 1 would misclassify one element." “We now simplify the complex expression from Step 12, which is $(48\times 15)-(320\div 8)+27$. The order of operations dictates multiplication and division before addition or subtraction. First compute $48\times 15=720$. Next compute $320\div 8=40$. Substituting these back into the expression gives 72040+27720-40+27. This reduction preserves equivalence while making the next step—performing the remaining subtraction and addition—more straightforward." Core Language Style Requirements Match the tone, sentence flow, and level of detail in the demonstration examples provided below. Write in a way that reads naturally to a human reasoner, not like a formal audit log. Self-Validation IMPORTANT: After each edge is generated, validate that whether it includes every step in direct_dependent_steps. If not, you need to re-generate the edge for the step. Notice that you need to check for all steps, not just a subset of steps. For each step, internally determine the applicable mathematical principle, identify dependencies, clarify evaluation or logic, and succinctly state both justification and role in the broader problem. Clearly state mathematical justifications, and follow the required language style. If a numeric evaluation is performed, confirm a brief sanity check is present. Verbosity edge should be detailed, allowing for expansion if the step is complex or multi-part. Stop Conditions Complete when every step has a well-justified edge field, following all style and content rules. Escalate for ambiguous or incomplete input only. Output Format Return only a edge field for each step populated as described. Do not include extra fields, wrappers, or commentary.
EXAMPLE JSON OUTPUT:
    "edges": [
        "step_id": 1,
        "edge": "We define ..."
        "step_id": 2,
        "edge": "Building on the definition of ..."
        "step_id": 3,
        "edge": "Using the expression for ..."
{JSON text of refined step-by-step answer with problem statement from Stage 1 and dependencies from Stage 2}
Appendix E Few-Shot Prompt for DAG-MATH Formatted CoT
[…]
  "steps": [
      "step_id": 36,
      "edge": "Using the confirmed digit values—
        K=0, L=5, M=3, and N=9, the sum K + L + M + N
        is expressed as 0 + 5 + 3 + 9. This step
        prepares the expression for final evaluation.",
      "direct_dependent_steps": [
        8,
        15,
        26,
        32,
        35
      "node": "The sum of the digits \$K + L + M + N\$
        is \$0 + 5 + 3 + 9\$."
2. The example below is bad since plural “Steps 8, 9, and 10” is used in edge instead of singular “Step 8, Step 9, and Step 10”.
  "steps": [
      "step_id": 11,
      "edge": "From Steps 8, 9, and 10, we have
        c(b - a) = 1 + 2k, b(c - a) = -3 + 6k,
        a(c - b) = -4 + 4k. To find relationships,
        assume a, b, c are such that differences are
        proportional. Let d = b - a, e = c - a,
        then c = a + e, b = a + d. Substitute into
        the equations.",
      "direct_dependent_steps": [
        8,
        9,
        10
      "node": "Define d = b - a and e = c - a. Then
      the equations become: (a + e) * d = 1 + 2k,
      (a + d) * e = -3 + 6k, a * (e - d) = -4 + 4k."
Problem: {Test problem statement} Solution:
Appendix F Additional Results for Evaluation
The results for final-answer accuracy (PASS@1) and empirical mathematical reasoning ability ($\widehat{\mathcal{R}}$) on three benchmarks across five LLMs are reported in Table˜2. The averaged graph-level statistics for BRUMO 2025 and HMMT 2025 are reported in Tables˜3 and 4, respectively. The overall trend is similar to the analysis in Section˜5. Additionally, we can obtain the following findings:
  • The change in graph-level statistics is monotonic in $\Delta$. The variations in the #nodes, #edges, density, and maximum out-degree from Correct to Perfect increase monotonically with the gap $\Delta$ between PASS@1 and $\widehat{\mathcal{R}}$. When $\Delta$ is small, the statistics of these two classes are nearly identical.
  • Graph-level statistics remain similar across the four classes when raw accuracy is low. In particular, when PASS@1 is low, their variations across classes are minimal.
Table 2: Final-answer accuracy (PASS@1) and empirical mathematical reasoning ability ($\widehat{\mathcal{R}}$) across three math benchmarks. Parentheses show the gap ($\Delta:=\text{PASS@1}-\widehat{\mathcal{R}}$).
Model AIME 2025 BRUMO 2025 HMMT 2025
PASS@1 $\widehat{\mathcal{R}}$ ($\Delta\downarrow$) PASS@1 $\widehat{\mathcal{R}}$ ($\Delta\downarrow$) PASS@1 $\widehat{\mathcal{R}}$ ($\Delta\downarrow$)
[…]
Std 10.3 2.1 12.2 2.5 11.0 0.9
Table 3: Averaged graph-level statistics of sampled DAGs across selected LLMs on BRUMO 2025.
Model Class #nodes #edges density ${\operatorname{d}}^{\max}_{\tt in}$ ${\operatorname{d}}^{\max}_{\tt out}$
Gemini-2.5-F All 26.6 39.8 13.0% 4.1 5.3
Incorrect 29.2 47.4 13.0% 4.8 6.2
Correct 25.1 35.9 13.1% 3.8 4.8
Perfect 20.5 26.1 14.1% 3.0 3.2
Gemini-2.5-F-L All 27.9 47.7 15.5% 3.8 8.3
Incorrect 33.2 61.5 14.5% 4.0 11.1
Correct 21.2 30.4 16.7% 3.4 4.8
Perfect 14.1 17.4 20.1% 2.6 2.9
GPT-4.1 All 14.9 18.1 19.4% 2.4 2.9
Incorrect 15.9 19.6 18.4% 2.5 3.0
Correct 13.0 15.1 21.4% 2.3 2.5
Perfect 12.0 13.8 23.0% 2.3 2.3
GPT-4.1-M All 17.1 24.1 18.7% 3.0 3.7
Incorrect 18.6 27.2 17.7% 3.1 4.1
Correct 14.3 18.3 20.6& 2.7 3.1
Perfect 13.1 16.6 22.1% 2.7 2.9
Qwen3-30B All 18.4 27.7 19.8% 3.4 4.7
Incorrect 21.7 34.4 17.8% 3.7 5.7
Correct 14.7 20.1 22.1% 3.1 3.6
Perfect 12.1 15.4 25.3% 2.8 3.0
Table 4: Averaged graph-level statistics of sampled DAGs across selected LLMs on HMMT 2025.
Model Class #nodes #edges density ${\operatorname{d}}^{\max}_{\tt in}$ ${\operatorname{d}}^{\max}_{\tt out}$
Gemini-2.5-F All 36.8 60.4 10.9% 4.7 7.9
Incorrect 36.6 61.4 11.4% 4.7 8.4
Correct 37.4 59.8 10.1% 4.9 7.1
Perfect 30.1 48.4 12.0% 5.9 6.1
Gemini-2.5-F-L All 35.8 63.0 14.0% 4.1 11.2
Incorrect 40.4 73.3 13.4% 4.1 13.1
Correct 26.5 42.7 15.2% 4.0 7.1
Perfect 16.8 25.9 20.6% 3.9 4.2
GPT-4.1 All 17.0 20.7 17.8% 2.5 3.2
Incorrect 16.9 20.8 17.8% 2.5 3.1
Correct 17.5 20.6 17.4% 2.6 3.5
Perfect 14.7 17.3 19.8% 2.5 2.7
GPT-4.1-M All 21.1 30.2 16.0% 3.0 4.2
Incorrect 21.1 30.2 15.9% 3.0 4.1
Correct 21.1 30.5 16.6% 3.0 4.4
Perfect 17.2 24.3 19.4% 2.9 3.6
Qwen3-30B All 22.8 34.6 16.1% 3.6 5.6
Incorrect 22.8 34.5 16.4% 3.6 5.7
Correct 22.8 35.0 15.2% 3.7 5.5
Perfect 18.7 28.5 18.8% 4.6 5.0